I need HELP!
I have a word problem, and i think i set it up right, but i dont think i solved it right, i got a negative speed?!?
Here is the problem: An airplane flew for 7 hours with a tail wind of 60 mph. The return flight against the same wind took 9 hours. Find the speed of the plane in still air.
this is how i set it up, please dont laugh:
(x-60)7=9(x+60)
then i distributed:
7x-420=9x+540
then put the variables on the same side:
7x-420=9x+540
-9x -9x
---------------
the 9x's cancel out, so it becomes:
-2x-420=540
then i get the "un"variables together, so it becomes:
-2x-420=540
+420+420
-------------
then the 420's cancel out, so it becomes:
-2x=960
then get the variable by itself by dividing:
-2x=960 or -2x/2 = 960/2
which then equals:
x=480 mph
now see this time i got something different, the 1st time i got -60 mph! PLEASE HELP0 ME!

Answer by venugopalramana(3286)

SEE BELOW
I have a word problem, and i think i set it up right, but i dont think i solved it right, i got a negative speed?!?
Here is the problem: An airplane flew for 7 hours with a tail wind of 60 mph.
LET SPEED OF PLANE IN STILL AIR =X
WIND SPEED =60
The return flight against the same wind took 9 hours. Find the speed of the plane in still air.
this is how i set it up, please dont laugh:
(x-60)7=9(x+60)...NO TAIL WIND HELPS THE FLIGHT SO ITS SPEED WITH TAIL WIND =X+60
AGAINST THE WIND THE SPEED IS X-60
SO EQN.IS
7(X+60)=9(X-60)
7X+420=9X-540
9X-7X=540+420=960
2X=960
X=480 MPH.

then i distributed:
7x-420=9x+540
then put the variables on the same side:
7x-420=9x+540
-9x -9x
---------------
the 9x's cancel out, so it becomes:
-2x-420=540
then i get the "un"variables together, so it becomes:
-2x-420=540
+420+420
-------------
then the 420's cancel out, so it becomes:
-2x=960
then get the variable by itself by dividing:
-2x=960 or -2x/2 = 960/2
which then equals:
x=480 mphNO YOU GOT X=-480....WHICH IS NOT CORRECT....
now see this time i got something different, the 1st time i got -60